# Structural Prognostic Event Modeling for Multimodal Cancer Survival Analysis

### v1

@@ G.1 Inference Runtime and Memory
We compare inference (where all patches are used) peak memory and wall-clock time against model performance on the COADREAD dataset. In this comparison, we set our model with slot number as 16 and iteration as 3. As shown in Figure 10(a), SlotSPE achieves the highest C-index while maintaining low runtime and memory usage. MCAT and MOTCAT achieve shorter runtime primarily because they omit self-attention computations within WSI patches; however, this design choice leads to a considerable drop in predictive performance. CMTA improves upon these baselines by employing Nyström attention (xiong2021nystromformer) as an approximation to full self-attention, but its capacity remains insufficient to capture the sparse, patient-specific prognostic signals. LD-CVAE attains the second-best performance, yet its reliance on a variational autoencoder introduces substantial memory and runtime overhead.
@@ H.1 Multivariable Survival Analysis
To assess whether SlotSPE provides prognostic value beyond clinical variables (age, sex, stage, and neoadjuvant treatment), which are empirically shown as crucial prognostic factors (xu2019deep; tomasello2017tumor; edge2010american; yu2022evaluating), we performed multivariable Cox regression (cox1972regression) using both clinical covariates and our model’s predicted risk scores. For each cohort, we merged the risk prediction of SlotSPE with these clinical variables on a per-patient basis. Importantly, the risk scores were obtained from the validation folds, ensuring that the model had no access to these samples during training. We then fitted Cox models across the five validation folds to quantify the incremental prognostic contribution of SlotSPE. In Table 10, we report the time-dependent AUC (heagerty2000time; heagerty2005survival), integrated brier score (IBS) (graf1999assessment; gerds2006consistent), and C-index for two configurations: (i) clinical-only, using only the clinical covariates, and (ii) clinical + model, where the predicted risk score of SlotSPE is included as an additional covariate. We also report the corresponding improvement C-index. Across all five cohorts, adding the risk score of SlotSPE to the clinical covariates consistently improves the C-index, with C-index ranging from 0.023 (KIRC) to 0.054 (LUAD). Time-dependent AUC likewise increases or remains comparable when including the model, while IBS is generally stable or reduced (e.g., substantial IBS reduction for KIRC and LUSC). These results indicate that SlotSPE captures complementary prognostic information that is not fully explained by routine clinical variables, showing the clinical utility potential to incorporate our model’s prediction as a clinically useful prognostic factor when combined with established clinical factors.
@@ I.2 Cohort-level Interpretability
In BRCA, for example, high-risk patients (bin1) are enriched in fatty acid and xenobiotic metabolism pathways, reflecting the metabolic reprogramming characteristic of aggressive tumors (pavlova2016emerging; rohrig2016multifaceted). In contrast, low-risk patients (bin4) exhibit stronger DNA repair and immune-related pathways, consistent with improved prognosis (nolan2017combined; denkert2018tumour). Such cohort-level analyses illustrate the potential of SlotSPE to uncover pathway–risk associations and generate hypotheses for further biological validation.

### v2

@@ G.1 Inference Runtime and Memory
We compare inference (where all patches are used) peak memory and wall-clock time against model performance on the COADREAD dataset. In this comparison, we set our model with slot number as 16 and iteration as 3. As shown in Figure 10(a), SlotSPE achieves the highest C-index while maintaining low runtime and memory usage. MCAT and MOTCAT achieve shorter runtime primarily because they omit self-attention computations within WSI patches; however, this design choice leads to a considerable drop in predictive performance. CMTA improves upon these baselines by employing Nyström attention (Xiong et al., 2021) as an approximation to full self-attention, but its capacity remains insufficient to capture the sparse, patient-specific prognostic signals. LD-CVAE attains the second-best performance, yet its reliance on a variational autoencoder introduces substantial memory and runtime overhead.
@@ H.1 Multivariable Survival Analysis
To assess whether SlotSPE provides prognostic value beyond clinical variables (age, sex, stage, and neoadjuvant treatment), which are empirically shown as crucial prognostic factors (Xu et al., 2019; Tomasello et al., 2017; Edge and Compton, 2010; Yu et al., 2022), we performed multivariable Cox regression (Cox, 1972) using both clinical covariates and our model’s predicted risk scores. For each cohort, we merged the risk prediction of SlotSPE with these clinical variables on a per-patient basis. Importantly, the risk scores were obtained from the validation folds, ensuring that the model had no access to these samples during training. We then fitted Cox models across the five validation folds to quantify the incremental prognostic contribution of SlotSPE. In Table 10, we report the time-dependent AUC (Heagerty et al., 2000; Heagerty and Zheng, 2005), integrated brier score (IBS) (Graf et al., 1999; Gerds and Schumacher, 2006), and C-index for two configurations: (i) clinical-only, using only the clinical covariates, and (ii) clinical + model, where the predicted risk score of SlotSPE is included as an additional covariate. We also report the corresponding improvement C-index. Across all five cohorts, adding the risk score of SlotSPE to the clinical covariates consistently improves the C-index, with C-index ranging from 0.023 (KIRC) to 0.054 (LUAD). Time-dependent AUC likewise increases or remains comparable when including the model, while IBS is generally stable or reduced (e.g., substantial IBS reduction for KIRC and LUSC). These results indicate that SlotSPE captures complementary prognostic information that is not fully explained by routine clinical variables, showing the clinical utility potential to incorporate our model’s prediction as a clinically useful prognostic factor when combined with established clinical factors.
@@ I.2 Cohort-level Interpretability
In BRCA, for example, high-risk patients (bin1) are enriched in fatty acid and xenobiotic metabolism pathways, reflecting the metabolic reprogramming characteristic of aggressive tumors (Pavlova and Thompson, 2016; Röhrig and Schulze, 2016). In contrast, low-risk patients (bin4) exhibit stronger DNA repair and immune-related pathways, consistent with improved prognosis (Nolan et al., 2017; Denkert et al., 2018). Such cohort-level analyses illustrate the potential of SlotSPE to uncover pathway–risk associations and generate hypotheses for further biological validation.